# Entanglement entropy between tangent balls in CFTD

### v1

@@ I Introduction
The purpose of this paper is to apply this universal approach in [1] to specific configurations and provide more verification. We consider the entanglement between two tangent balls, as shown in the left panel of Figure 2. We also compute the entanglement entropy of a ball and a half-space, obtained as the special case where the radius of one ball is taken to infinity, as shown in the left panel of Figure 3. In CFT2, this configuration corresponds to entanglement in boundary conformal field theory (BCFT) between an interval and the negative half-axis, as depicted in the right panel of Figure 3.
@@
The structure of the paper is as follows: In Section II, we apply the universal method in [1] to the tangent case and further compute the entanglement entropy between a ball and a half-space, which arises as a special case of the tangent configuration. In Section III, we calculate the same entanglement in CFT2 by subtraction method [7] and show that two-dimensional solid torus approach coincides with subtraction method.
@@ II Entanglement entropy in the tangent configuration and its half-space limit
In this section, we first apply the universal approach to compute the entanglement entropy in the tangent configuration. We then consider the half-space limit by taking the radius of one ball to infinity and evaluate the corresponding entanglement entropy.
@@ II.2 Half-space limit
For another simpler case, the entanglement of two half-space with two cutoffs $l_{1}=\epsilon\to 0$ and $l_{2}=2\xi\to\infty$, we can easily derive that the entanglement entropy is:
@@ III A specific example of computation in CFT2
which precisely corresponds to the previous results of entanglement entropy for an interval with length 2R in the half-space.
In fact, this result can be obtained by subtraction method in a similar way. As [6, 7, 8] has illustrated, for two disjoint intervals in CFT2 at the time slice $\tau=0$, we can remove two discs from the Euclidean path integral region to construct a pure state $\psi_{AB}$. The doubly connected region can then be conformally mapped onto an annulus with width $W=\log\frac{r_{2}}{r_{1}}$ and two boundary states $|a,b\rangle$, as illustrated in the right panel of Figure 6. The entanglement entropy in the annulus is defined in terms of the Rényi entropy:
@@
Figure 7: This picture shows the half-space limit case in $\mathrm{CFT}_{2}$. Region A is the negative half-axis ($\infty,\xi_{2}$), while region B is the finite interval terminating on the boundary ($\xi_{3},\xi_{4}$). The two intervals are separated by a small distance $2\epsilon=\xi_{3}-\xi_{2}$, which serves as a UV regulator.
To reduce the general case to the half-space BCFT, we consider the negative half-axis (region A) and an interval (region B) by setting $\xi_{1}=$-$\infty$, $\xi_{2}=-2R-\epsilon$, $\xi_{3}=-2R+\epsilon$, and $\xi_{4}=\epsilon\to 0$, as shown in Figure 7. The cross ratio is given by $\eta=\frac{R}{\epsilon}$. Substituting Eq. (23) into the expression for $S_{\text{vN}}(A:B)$, we obtain the following result:
@@
which is in agreement with equation (19), the half-space entanglement entropy with boundaries. In the AdS/CFT correspondence, the entanglement entropy of the half-space corresponds to half of the geodesic cutoff [12].

### v2

@@ I Introduction
The purpose of this paper is to apply this universal approach in [1] to specific configurations and provide more verification. We first consider the entanglement between two tangent balls, as shown in the left panel of Figure 2. We then compute the entanglement entropy of a ball and a half-space: in the tangent case this is obtained as the special limit where the radius of one ball is sent to infinity, as shown in the left panel of Figure 3, while in the nonadjacent configuration the ball is separated from the half-space by a finite distance. In CFT2, this configuration corresponds to entanglement in boundary conformal field theory (BCFT) between an interval and the negative half-axis, as depicted in the right panel of Figure 3.
@@
The structure of the paper is as follows: In Section II, we apply the universal method in [1] to the tangent case and further compute the entanglement entropy between a ball and a half-space, which arises as a special case of the tangent configuration. We also compute the entanglement entropy for a configuration in which the ball and the half-space are spatially separated. In Section III, we calculate the same entanglement in CFT2 by subtraction method [7] and show that two-dimensional solid torus approach coincides with subtraction method.
@@ II Entanglement entropy in the tangent configuration and its half-space limit
In this section, we first apply the universal approach to compute the entanglement entropy in the tangent configuration. We then consider the half-space limit by taking the radius of one ball to infinity and evaluate the corresponding entanglement entropy. Finally, we consider a nonadjacent configuration in which the ball is separated from the half-space by a finite distance and evaluate the corresponding entanglement entropy.
@@ II.2 Half-space limit
For another simpler case, the entanglement of two half-spaces with two cutoffs $l_{1}=\epsilon\to 0$ and $l_{2}=2\xi\to\infty$, we can easily derive that the entanglement entropy is:
@@
II.3 Nonadjacent half-space limit
For completeness, we also discuss the half-space limit in the nonadjacent configuration, shown in the left panel of Figure 5. In this setup, the ball and the half-space are separated by a finite distance $d$. The inversive product is
$\varrho=\frac{\left|\vec{n}\cdot\vec{x}\right|}{r}=\frac{R+d}{R}.$ (18)
Thus the entanglement entropy takes the form
$S_{\mathrm{disj}}(A:B)=-4\mathcal{E}_{\text{vac}}\pi^{\frac{D}{2}}\frac{\Gamma\left(\frac{D}{2}\right)}{\Gamma(D)}\left(\frac{4R}{d}+2\sqrt{\frac{4R^{2}}{d^{2}}+\frac{2R}{d}}\right)^{D-1}{}_{2}F_{1}\left(D-1,\frac{D}{2};D;-\frac{4R}{d}-2\sqrt{\frac{4R^{2}}{d^{2}}+\frac{2R}{d}}\right).$ (19)
We also record the explicit expressions in several dimensions:
$D=2$ $:\quad 8\pi\mathcal{E}_{2}\log\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right),$
$D=3$ $:\quad 4\pi^{2}\mathcal{E}_{3}\frac{\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}-1\right)^{2}}{\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}},$
$D=4$ $:\quad 2\pi^{2}\mathcal{E}_{4}\left[\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{2}-\frac{1}{\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{2}}-4\log\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)\right],$
$D=5$ $:\quad\frac{\pi^{3}\mathcal{E}_{5}}{3}\left[\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{3}+\frac{1}{\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{3}}\right$.
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\left.-9\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}+\frac{1}{\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}}\right)+16\right],$
$D=6$ $:\quad\frac{\pi^{3}\mathcal{E}_{6}}{6}\left[\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{4}-\frac{1}{\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{4}}\right$.
$\left.-8\left(\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{2}-\frac{1}{\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)^{2}}\right)+24\log\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)\right].$ (20)
From this expression, one can readily see that the entanglement entropy in the nonadjacent configuration is finite. In particular, for D=2, the expression is
$S=8\pi\mathcal{E}_{2}\log\left(\sqrt{\frac{2R}{d}}+\sqrt{1+\frac{2R}{d}}\right)=4\pi\mathcal{E}_{2}\log\left(1+\frac{4R}{d}+2\sqrt{\frac{2R}{d}\left(\frac{2R}{d}+1\right)}\right).$ (21)
This expression precisely matches the entanglement entropy of a quantum state with an interval of length $2R$ and a semi-infinite line that is separated from it by a distance $d$, as will be illustrated in the next section. By treating the finite separation $d$ as an infinitesimal UV regulator and sending $d\to 0$, the nonadjacent result smoothly reduces to the tangent one.
@@ III A specific example of computation in CFT2
which precisely corresponds to the previous results of entanglement entropy for an interval with length $2R$ in the half-space. Moreover, the nonadjacent entanglement entropy in equation (21) turns to
$S=\frac{c}{6}\log\left(1+\frac{4R}{d}+2\sqrt{\frac{2R}{d}\left(\frac{2R}{d}+1\right)}\right).$ (24)
This result also precisely matches the entanglement entropy for an interval with length $2R$ and the negative half-axis separated by a distance $d$, as shown in the right panel of Figure 7.
In fact, these results can be obtained by subtraction method in a similar way. As [6, 7, 8] have illustrated, for two disjoint intervals in CFT2 at the time slice $\tau=0$, we can remove two discs from the Euclidean path integral region to construct a pure state $\psi_{AB}$. The doubly connected region can then be conformally mapped onto an annulus with width $W=\log\frac{r_{2}}{r_{1}}$ and two boundary states $|a,b\rangle$, as illustrated in the right panel of Figure 6. The entanglement entropy in the annulus is defined in terms of the Rényi entropy:
@@
Figure 7: This picture shows the half-space limit case in $\mathrm{CFT}_{2}$. In the left panel, region A is the negative half-axis ($\infty,\xi_{2}$), while region B is the finite interval terminating on the boundary ($\xi_{3},\xi_{4}$). The two intervals are separated by a small distance $2\epsilon=\xi_{3}-\xi_{2}$, which serves as a UV regulator. In the right panel, region A and region B are separated by a finite distance $d$.
To reduce the general case to the half-space BCFT, we consider the negative half-axis (region A) and an interval (region B) by setting $\xi_{1}=$-$\infty$, $\xi_{2}=-2R-\epsilon$, $\xi_{3}=-2R+\epsilon$, and $\xi_{4}=-\epsilon\to 0$, as shown in the left panel of Figure 7. The cross ratio is given by $\eta=\frac{R}{\epsilon}$. Substituting equation (28) into the expression for $S_{\text{vN}}(A:B)$, we obtain the following result:
@@
which is in agreement with equation (23), the half-space entanglement entropy with boundaries. In the AdS/CFT correspondence, the entanglement entropy of the half-space corresponds to half of the geodesic cutoff [12]. For the nonadjacent configuration, we take $\xi_{1}^{\prime}=$-$\infty$, $\xi_{2}^{\prime}=-2R-d$, $\xi_{3}^{\prime}=-2R$, and $\xi_{4}^{\prime}=-\epsilon\to 0$, as depicted in the right panel of Figure 7. The corresponding cross ratio is $\eta^{\prime}=\frac{2R}{d}$ and the entanglement entropy becomes
$S_{\text{vN}}(A:B)=\frac{c}{6}\log\left(1+\frac{4R}{d}+2\sqrt{\frac{2R}{d}\left(\frac{2R}{d}+1\right)}\right),$ (30)
which is again in agreement with equation (24).